### v2

# Effect of ultralight dark matter on compact binary mergers

Abstract
The growing catalogue of gravitational wave events enables a statistical analysis of compact binary mergers, typically quantified by the merger rate density. This quantity can be influenced by ambient factors, following which, in this work we have investigated the impact of dark matter environment on the merger statistics. We construct a baseline astrophysical model of compact binary mergers and extend it by incorporating a model of ultra light dark matter, which affects the orbital evolution of binaries through accretion and dynamical friction. Our analysis of the merged population of binary progenitors demonstrates that, compared to the baseline model, ULDM can significantly alter the merger statistics when its ambient density becomes larger than $10^{4}\textrm{GeV}/\textrm{cm}^{3}$. A comparison with the gravitational wave data from the GWTC-3 catalogue provides insight into potential observational signatures of the ULDM in merger events, leading to possible constraints on the existence and density of dark matter distribution in galaxies.
I Introduction
Understanding the nature of the compact objects, e.g., black holes (BHs) [12, 9], neutron stars (NSs) [59, 2] and exotic compact objects (ECOs) [21], that exist in our universe and assessing their environments has been one of the key research areas of gravitational physics. This will not only provide information about the properties of these compact objects and strong gravity regime that these objects live in, but also about their environments [10, 21]. The upshot being, we will be able to confidently comment, whether general relativity is the theory of gravity describing our universe even in the strong field regime, as well as prove the existence of objects, that are more compact than NSs, but less compact than BHs [27, 52, 13, 20]. These will revolutionize the current understanding of our universe in many possible ways .
However, the significance of studying compact object environments extends beyond these aspects. It can also reveal crucial information about a variety of astrophysical phenomena, including the presence of a third body—such as exoplanets or stars—orbiting these compact objects [57, 51]. Additionally, such investigations can shed light on the existence and distribution of dark matter (DM) [46, 26, 14, 19, 22, 15]. In this work, we focus specifically on exploring the presence and properties of DM based on current observational data.
As far as assessing the physics on and around compact objects [^1] Throughout this work, by compact objects, we will mean objects, whose compactness is larger or equivalent to that of a typical NS. are considered, the black hole shadow [50] and the gravitational wave (GW) observations are most effective. While, observations involving black hole shadow are severely limited, in the sense that only two BHs, the M87* and the Sgr A* has been pictured so far [6, 7], numerous GWs from compact binary coalescence have already been detected through the LIGO-VIRGO-KAGRA (LVK) collaboration [1, 2] and the detection rate is expected to grow exponentially with future generations of GW detectors [47, 53]. Thus, we will focus on understanding the properties of DM distribution around compact objects through the GW observations.
[…]
One way to circumvent this difficulty, is to assume that DM form a condensate. This is where, the models of Ultra-Light DM (ULDM) play a pivotal role [29] and can alleviate some of the challenges faced by the CDM scenario. In this formulation, DM is thought to be composed of light bosons, which forms a condensate at sub-halo length scales, while smoothly joins with CDM at larger distances, without violating any observational constraints [29]. Moreover, such light bosons have also been shown to arise from simple extensions of the standard model for particle physics, e.g., models involving axions [31]. All of these make ULDM models a viable extension of the CDM at the galactic scale.
At the same time, as elaborated above, mergers of compact objects have now become a regular occurrence since their discovery in 2015 by the LVK detectors, which have now detected $\mathcal{O}(100)$ merger events. This presents an interesting possibility to check the existence of ULDM at galactic scales through GW observations. The basics are as follows, in the presence of an environment consisting of ULDM, the binary objects are moving around each other not in vacuum, but in a medium composed of ULDM. Consequently, both the objects in the binary will experience a viscous drag due to the presence of ULDM medium, which can also accrete into the individual objects. The calculation of the viscous drag force in the relativistic context was shown in [58] to be a straight generalisation of the Newtonian dynamical friction formula derived by Chandrasekhar [24]. It turns out that both the dynamical friction, as well as accretion, as described above, alters the binary inspiral dynamics and leave their signature on the binary phasing. This concept was used [18] to constrain the ambient density of baryons from specific merger events in GWTC2 [3]. More recently another work [49] had also used ULDM interactions to distinguish different surface properties of the compact objects, though using extreme-mass-ratio inspirals.
In this work, we start with a simple implication of the previous literature, namely that an environment is likely to alter the merger timescales for a binary system. Given this alteration, we investigate if the ULDM environment can leave its biasing upon the statistics of compact binary mergers. In other words, we want to understand how big or small the effect of ULDM dynamical friction and/or accretion is upon the number of mergers happening per unit redshift per unit observer time per unit co-moving volume, the so-called ‘astrophysical merger rate density’. Due to the inherent astrophysical uncertainty in the star formation rate, and its modelling, the results presented here will not be conclusive, rather will provide a proof-of-principle that DMs and their properties can be understood using merger rate density. Further astrophysical refinements will make our conclusions more robust, which we will come back to in a subsequent publication. Our paper is organised as follows: in Section II, we review some of the key previous results which are necessary for our subsequent calculations involving the biasing effects expected for mergers in an ULDM environment in Section III. We take help of toy problems to demonstrate the calculations for individual binaries in Section III.1 and populations Section III.2. We present our results for the astrophysical merger rate of compact objects in presence of an ULDM environment in Section IV. We finally present our conclusions in Section V.
Conventions — Throughout this paper we will use geometric units, namely $G=c=1$. Note that we do not use $\hbar=1$. Also, we work in four spacetime dimensions, with mostly positive signature convention, such that the flat metric becomes $\eta_{\mu\nu}=\textrm{diag.}(-1,+1,+1,+1)$.
II Binary mergers in an ultra-light dark matter environment
In this section, we will present a brief review of the framework and the implications of the ULDM environment on the merger rate of binary compact objects. We start by reviewing the computation of accretion and dynamical friction experienced by a compact object in this environment, following [58, 49]. Our implementation of these effects into the dynamics of binary mergers and their rate will follow from analogous works involving visible matter environments [18, 54].
II.1 Framework
We consider the ULDM to be represented by a light scalar field, which satisfies the massive Klein-Gordon equation
$\Box\Phi-\mu^{2}\Phi=0\,$, (1)
where $m_{\Phi}=\hbar\mu$ is the mass of the scalar field. The box operator in Eq. (1) is taken to be a non-trivial background given by the geometry of the massive compact object. If the compact object is non-rotating, the geometry will be that of Schwarzschild, while for rotating compact object the geometry will be given by Kerr metric. Then, following [49], we look for solutions for the scalar field at the rest frame of the compact object, which at large distance from the compact object, behave as plane waves,
$\Phi(t,r)\sim\sqrt{\frac{\rho_{0}}{\mu}}\,e^{i(\omega t-k_{\infty}r)}\,.$ (2)
Here, $\rho_{0}$ is the homogenous rest mass density of the scalar field, and $k_{\infty}=\sqrt{\omega^{2}-\mu^{2}}$. In general, the density $\rho_{0}$ of the ULDM, appearing above, will be dependent on the distance from the galactic halo, however, we will assume it to be homogeneous, since our interest is to observe if there is at all any effect of DM on the binary merger rate. Subsequently, one can refine the analysis further, by introducing more realistic density profiles for ULDM. Note that for $\omega>\mu$, the scalar particle scatters off the compact object, while for $\omega<\mu$, the scalar particles form quasi-bound states. In this computation, we will assume that $\omega>\mu$, along with $M\omega\ll 1$, which is the prime reason for studying the ultra-light sector, so that the above conditions are trivially satisfied. The relative velocity of the BH with respect to the DM particles is given by $v^{i}=-k^{i}_{\infty}/\omega$, with $|k^{i}_{\infty}|=k_{\infty}$. Therefore, it follows that, $\omega=\gamma\mu$ and $k^{i}_{\infty}=-\gamma\mu v^{i}$, with $\gamma=1/\sqrt{1-v^{2}}$ is the Lorentz factor.
The computation of the rate of accretion and dynamical friction follows the below steps — (a) one solves for the radial part of Eq. (1) in the Kerr background, in the asymptotic and the near-horizon regime; (b) for the near-horizon solution, one imposes reflecting boundary conditions, when these binary objects are stars, and impose purely ingoing boundary conditions for BHs; (c) the asymptotic limit of the near-horizon solution is matched with the near-zone limit of the asymptotic solution; (d) this matching provides ratio of reflected and incident amplitudes of the scalar field, which in turn provides the accretion rate and force due to dynamical friction. In the frame of the object moving through DM distribution, these read [49, 58],
$\dot{M}$ $=\int_{\infty}r^{2}T_{tr}d\Omega_{2}=\rho A_{\rm h}\left(\frac{e^{\pi\eta}\pi\eta}{\sinh(\pi\eta)}\right)\left(\frac{1-R}{1+R}\right)~,$ (3)
$F^{i}$ $=\int_{\infty}r^{2}T^{i}_{r}d\Omega_{2}=-\frac{4\pi\rho M^{2}}{v^{2}}(1+v^{2})^{2}\mathcal{D}\zeta^{i}-\dot{M}v^{i}~.$ (4)
The quantity $R$, appearing above, in the expression for $\dot{M}$, is unity for stars, and vanishes for BHs. Thus the above expression for accretion is relevant for BHs alone. Here, $T_{\mu\nu}$ is the energy-momentum tensor of the scalar field, $\zeta^{i}$ is the direction from which the scalar waves impinges on the BH at rest, $\eta=-\{\gamma\mu M(1+v^{2})/v\}$, $\rho=\gamma^{2}\rho_{0}$ and $\mathcal{D}=\ln\Lambda-\textrm{Re}[\Psi(1-i\eta)]$. The quantity $\Psi(x)$ is the di-gamma function and $\Lambda=\gamma v\mu b_{\rm max}$, with $b_{\rm max}$ being the maximum impact parameter. Note that $\gamma v\mu$ is the inverse of the de-Broglie wavelength associated with the DM particles. Finally, the quantity $A_{\rm h}=4\pi(r_{+}^{2}+a^{2})$ is the horizon area, with $r_{+}=M+\sqrt{M^{2}-a^{2}}$, being the location of the event horizon, and $J=aM$ is the angular momentum of the BH. In the expression for force experienced by the compact object due to scattering from ULDM, the first term represents the gravitational drag from the surrounding ULDM field, and the second term accounts for momentum loss due to accretion.
The above analysis was performed in the frame of the compact objects, while the observation is being made at asymptotic infinity. Since the DM is static in the observer’s frame of reference, we need to make a transformation from the frame of the BH to the frame of the observer. This is simply achieved by the following transformations: $\dot{M}^{\prime}=(\dot{M}/\gamma)$ and $F^{\prime}_{i}=F_{i}+\dot{M}v_{i}$. Then, it turns out (see [49] for the detailed calculation) that the accretion rate of an object of mass $M$ and angular momentum $J$ moving through the ULDM medium with an orbital speed $v$, in the rest frame of the DM, is given by,
$\dot{M}^{\prime}=\hbar n\omega A_{+}\left(\frac{e^{-\pi\eta}\pi\eta}{\sinh(\pi\eta)}\right)\left(\frac{1-R}{1+R}\right),$ (5)
where, $n=(\rho/m_{\Phi})$ is the number density of the ULDM particles, and we assume complete absorption if the object is a BH ($R=0$) and complete transmission in all other cases ($R=1$). Similarly, the dynamical friction force acting on the object in the rest frame of the DM is given by
$\vec{F}^{\prime}$ $=-\frac{4\pi\rho M^{2}}{v^{2}}(1+v^{2})^{2}\mathcal{D}\vec{\zeta}=-\frac{4\pi\hbar n\eta^{2}}{\mu}\Delta\Psi\vec{\zeta}\,$, (6)
where, $\mathcal{D}=\Delta\Psi=\mathrm{Re}[\Psi(1+\ell_{m}+i\eta)-\Psi(1+i\eta)]$ with $\Psi$ being the digamma function and $\ell_{m}$ the cut-off in the angular harmonic corresponding to the maximum impact parameter of the object $b_{\rm max}$. As evident, the dynamical friction is independent of the details of the nature of the compact object, while accretion is dependent on the compact object. We will show the effect of accretion on binary merger rate in Section II.2. Multiplying the dynamical friction force by the relative speed $v$ between the compact objects, we can finally deduce the power dissipation through the dynamical friction as
$P_{\mathrm{df}}=\vec{F}^{\prime}\cdot\vec{v}=-\frac{4\pi\hbar}{\mu}nv\eta^{2}\Delta\Psi~.$ (7)
This dissipative term will add up to the quadrupolar radiation emitted from the binary inspiral, leading to faster energy loss from the system. This results into faster inspiral, and hence reduces the merger time, having impact on the rate of binary merger. We will implement this for binary orbits in the subsequent section.
II.2 Implementation in binary inspiral
Following the outlines of Section II.1, we note that the motion of binary objects through the DM environment creates an extra dissipation force, which in the scalar’s rest field gives rise to an additional loss of power, given by $P_{\mathrm{df}}$. In the case of a binary orbit, the consequence being, the loss of the total orbital energy of the binary. In addition to this dissipation due to DM, there is also a dissipation from the emission of GWs, which at leading order is given by [60],
$P_{\mathrm{gw}}=\frac{32}{5}\frac{m_{\rm r}^{2}M_{\rm T}^{3}}{a^{5}}\,$, (8)
where, $m_{\rm r}=M_{1}M_{2}/(M_{1}+M_{2})$ and $M_{\rm T}=M_{1}+M_{2}$ are the reduced mass and the total mass of the particular binary, respectively, with individual masses being $M_{1}$ and $M_{2}$. The quantity $a$ is the binary separation and is related to the orbital velocity $v$, as well as the orbital frequency $\Omega$ of the reduced mass system by $v^{2}=(M_{\rm T}/a)=(M_{\rm T}\Omega)^{2/3}$. Note that the frequency of the emitted GW is twice of the orbital frequency. Following the implementation of [54], the equation for orbital phasing of GW becomes
$\frac{d\phi}{dt}$ $=\Omega\,$,
$\frac{dE_{\rm orb}}{dt}$ $=-P_{\mathrm{gw}}-P_{\mathrm{df}}\,$, (9)
where, the orbital energy $E_{\rm orb}$ is given at the leading order by $E_{\rm orb}=-(1/2)m_{\rm r}v^{2}$. We can then substitute for $E_{\rm orb}$ along with the expressions of $P_{\mathrm{df}}$ and $P_{\mathrm{gw}}$ from Eq. (7) and Eq. (8) to obtain,
$\frac{1}{2}\frac{d}{dt}\left(\frac{m_{\rm r}M_{\rm T}}{a}\right)=-\frac{32}{5}\frac{m_{\rm r}^{2}M_{\rm T}^{3}}{a^{5}}-\frac{4\pi\hbar n}{\mu}\sum_{i=1}^{2}\eta^{2}_{i}v_{i}\Delta\Psi_{i}\,$, (10)
where the sum over $i$ represents a summation over the components of the binary. Then it is easy to see from the motion of the reduced mass that $v_{i}=(m_{i}/M_{\rm T})v=(m_{i}/M_{\rm T})\sqrt{M_{\rm T}/a}$, where $i,j$ are just distinct labels for the binary components and not spatial indices. We should also note that the time differentiation operator on the left hand of Eq. (10) will also hit both $m_{\rm r}$ and $M_{\rm T}$, where the corresponding expressions have to be substituted from Eq. (5). In this way we see that we end up with a differential equation in terms of the binary’s orbital separation $a$. The numerical evolution of Eq. (10) determines the dynamics of the binary in question, which will be different from the vacuum evolution, due to the presence of dynamical friction.
III Dark matter environment and merger statistics
Here, we will present our results regarding merger statistics of binary evolution in the environment of ULDM and shall quantify the effects of ULDM on compact binary mergers. We start by discussing the effects of environment on individual binaries and then focus on the effect on binary population and their merger rate.
III.1 Individual Binaries
We show the effect of the presence of ULDM on the merger timescale of individual binaries in this section. For that we start with a chosen mass configuration of binary main sequence stars as progenitors and track their merger status by evolving Eq. (II.2) to obtain the binary separation $a$ with time. To do this consistently, we require a proper transformation function when a progenitor star reaches its stellar endpoint. We implement this by a simple mass-transfer function [32] to determine the mass of the evolutionary remnant of the main sequence star. In reality, such endpoints of massive main-sequence stars are usually accompanied by novas, which are highly model dependent and is likely to add a further layer of uncertainty to our baseline model, but will not significantly contribute to an understanding of the effects of ULDM on binary merger. Nevertheless, we need to implement a model to facilitate mergers and not disrupt them. To do this, we adopt a method similar to the evolution method previously implemented in [23] with the notable exception that our method does not implement a constant ejecta velocity to model the kick imparted to the stellar remnants. Rather, we let the ejecta velocity to depend on the initial masses of the stellar progenitors and also on the associated mass loss. In this case, the kick imparted to the remnant is likely to cause a sudden change to the angular momentum, depending on various parameters of the system. The change in angular momentum also causes a change in the angular velocity of the remnants. These changes can be determined through the following equations obtained from $L\coloneqq m_{\rm r}a^{2}\Omega$ and Kepler’s third law $\Omega^{2}a^{3}=M_{\rm T}$.
$\left(\frac{\delta\Omega}{\Omega}\right)$ $=-3\left(\frac{\delta L}{L}\right)+3\left(\frac{\delta m_{\rm r}}{m_{\rm r}}\right)+2\left(\frac{\delta M_{\rm T}}{M_{\rm T}}\right)\,.$ (11)
Here $\delta x$ represents the change in a quantity $x$ before and after the stellar endpoint. From the mass-transfer function implementation of [32] the final mass is less than the initial mass because a nova usually involves ejection of mass. Therefore, it follows that $(\delta M_{\rm T},\delta m_{\rm r})<0$. Using these results in Eq. (11), we notice that only if $\delta L<0$, one might have $\delta\Omega>0$. Then to connect with the binary separation at the stellar endpoint, we use Kepler’s third law again to obtain,
$3(\delta a/a)=(\delta M_{\rm T}/M_{\rm T})-2(\delta\Omega/\Omega)$ (12)
This suggests that whenever $(\delta\Omega/\Omega)\geq 0$, the binary separation always decreases at stellar endpoint. Therefore we work with precisely those events where $(\delta\Omega/\Omega)\geq 0$ is obeyed at mass transfer, which in turn ensure that the binary separation never increase across the stellar endpoint. We note that the condition $\delta L<0$ implies that the angular momentum induced by the ejecta is opposite to that of the orbital angular momentum. On the other hand, the condition $\delta L\geq 0$ increases the binary separation and can lead to potential disruption of the system [11]. Hence on average, the number of binaries whose separation between the individual components decrease will be reduced by some factor. Luckily, this decrease in the number of such events does not affect our results because we rescale the merger events so that the local merger rate density matches observations (see Sec  IV).
Further, from the above expressions it follows that $|\delta a/a|\geq(1/3)|\delta M_{\rm T}/M_{\rm T}|$. From this, we finally make our choice of $\delta a$ across stellar endpoint as
$\delta a=\left(\frac{\delta M_{\rm T}}{M_{\rm T}}\right)a\,.$ (13)
Note that, the above relation ensures that $\delta M_{\rm T}<0$, implies $\delta a<0$, since both $a$ and $M_{\rm T}$ are positive definite quantities. This completes the formulation of our baseline model, which is constructed in the absence of any ULDM environment.
Before moving further we note that the overall effect of the ULDM environment is through accretion and dynamical friction. As already emphasized, Eq. (6) depicts that the drag term acts universally upon any object, but the accretion term is dependent upon the nature of the surface reflectivity, which will vanish for a BH, while will be unity for a star. Thus the evolution of the binary separation will be different for stars and BHs. With this specific model, we consider the following example involving a binary progenitor configuration of total mass and mass-ratio being $M_{\rm T}^{\rm P}=45M_{\odot}$ and, $q^{\rm P}=0.8$, respectively, for two different starting redshift and for two cases with and without ULDM environment. Quantities related to the progenitor stars will be denoted with a superscript P. At this point we remark that further simplicity is maintained in our baseline model by not considering properties like orbital eccentricity and/or BH spins. Under these simplifying assumptions, any difference in the final binary configuration between the baseline model and the model with ULDM can be thought of as being solely due to the presence of the ULDM environment.
Figure 1: The figure depicts the evolution of the binary separation parameter $a$ (in units of solar radius $R_{\odot}$) of a given progenitor as a function of the redshift $z$. We assume that the progenitors start out at $z=0.2$ and consists of binary main sequence stars with $M_{\rm T}^{\rm P}=45M_{\odot}$ and $q^{\rm P}=0.8$. The left most region of the plot corresponds to $z\sim 0.2$, while the rightmost region is $z\sim 0$. The initial colour shading (brown) corresponds to the star-star phase, the next colour shading (green) corresponds to the star-compact object phase and the final colour shading refers to the situation when both are compact objects. We show our results with ($\rho>0$) and without ($\rho=0$) the presence of ULDM. The units of DM density $\rho$ are in $\textrm{GeV}/\textrm{cm}^{3}$. See text for more details.
The results for the evolution of the orbital separation of the binary system, with progenitors having $M_{\rm T}^{\rm P}=45M_{\odot}$ and, $q^{\rm P}=0.8$, have been presented in Fig. 1, starting at redshift $z=0.2$. The evolutionary phases involve the binary components to be either stars or compact objects, e.g., NSs and BHs. Starting from higher redshift values (left hand side of Fig. 1), the different shadings correspond to the binary in its different phases starting from the star-star phase (brown shade), the compact object-star phase (green phase) and finally when both are compact objects (the crimson shade). As per our elaboration of the baseline model, we notice two discontinuous jumps in the binary separation $a$ across the two endpoints of the main-sequence stars in the binary, following the implementation in Eq. (13). These endpoints of the stars are determined through the mass-luminosity relation of main sequence stars [55], which allows us to compute a simple scaling relation for the lifespan $\tau_{\rm star}$ of a star with its progenitor mass $M^{\rm P}$, namely $\tau_{\rm star}\propto M^{-2.5}$. Dividing the corresponding quantities for the Sun, and keeping in mind that the estimated life span of Sun is $\tau_{\odot}\approx 10$ Gyr, we obtain the typical lifespan of a star in the binary to be,
$\tau^{i}_{\rm star}(\textrm{in}~\mathrm{Gyr})=\frac{10}{({M^{\rm P}_{i})}^{2.5}}\,.$ (14)
Here, $i$ refers to the binary components. We note that owing to the source starting from $z=0.2$, and the previous relation for the lifespan of main-sequence stars, the binary remains unmerged at $z=0$ even in the presence of ULDM. However, even the unmerged binaries will carry a signature of the DM environment in their present orbital separation $a_{0}\equiv a(z=0)$, which for the ULDM case is almost half its value for the vacuum configuration (see the horizontal dashed lines in Fig. 1).
We would like to point out that unmerged binaries are effectively binaries in their early inspiral, which cannot be observed by the ground based GW detectors. But, un-merged binaries will show up as nearly monochromatic GWs in the space based detectors, e.g., LISA [37] and DECIGO [38]. Given that the orbital separation $\mathrm{a}$ is highly dependent on the desnity of DM particles in the environment, independent constraints on the ULDM can be obtained from the observation of low-frequency GW signals by LISA [28].
Figure 2: Same configuration as in Fig. 1, except for the fact that the progenitors start out at $z=0.8$. In this case, the objects merge at a higher redshift (z=0.4) in the presence of ULDM environment, while in vacuum they merge much later, close to $z=0$. The ‘0’ tick on the x-axis is just to guide the eye.
Finally, we present the effect of ULDM on a similar binary configuration as before, except for the fact that the binary progenitors now start at a higher value of redshift, namely at $z=0.8$. This scenario is presented in Fig. 2. In contrast to the previous case, we see that merger happens both in the presence as well as in the absence of ULDM. However, as evident from Fig. 2, the presence of ULDM is seen to make the merger faster, thereby increasing the merger redshift from $z=0.07$ to $z=0.4$. This behaviour is expected, and in consonance with the results that we obtained earlier. In short, the dissipation from the ULDM environment adds to the energy loss of the binary, over and above the energy loss from GWs, and consequently we are left with faster merger timescales and/or smaller remnant separation.
III.2 Evolution of a binary population of progenitors from a given redshift
We have discussed in the previous section how the ULDM environment affects a pair of compact objects undergoing a binary inspiral. In this section we will present the corresponding result for a population of binary progenitors. In doing so, the individual pair of objects can now be thought of as building blocks of a population. To characterise the statistics of merger populations, we need to answer to the question, what happens when many of such individual binaries evolve under a ULDM environment simultaneously. We will answer this question by first considering a toy problem, namely, we will determine the final configuration of a population of binary main-sequence star systems, all of which began at the same redshift. All of these binary systems, during their evolution, undergo accretion from the ULDM environment (if they gets converted to BHs) as well as dynamical friction. In addition to these, they will lose energy by emission of GW. The rate of dissipation of the orbital energy, therefore, depends on all the binary parameters (namely masses, spins and eccentricity) as well as on the environment the binary system lives in. Hence, it is clear that the orbital evolution of the binaries will non-trivially depend on the progenitors mass distribution $\{M_{i}^{\rm P}\}$ (the so-called initial mass function or, IMF), the spin distribution $\{\chi_{i}^{\rm P}\}$ ($\chi\equiv(J/M^{2})$, where $J$ is the angular momentum) and the eccentricity distribution $\{e^{\rm P}\}$ of the population of the progenitor stars and their evolved remnants. The IMF is for the most part unknown, but is expected to be intricately linked to astrophysical factors like stellar metallicity and accretion. The only estimates of the IMF comes from observations of stars in our own galaxy [42] and is usually taken to follow a broken power-law distribution with less massive stars being more probable. Without observation of individual stars, very little can be known about the nature of the IMF [41, 36] in a model independent way at the redshifts of star formation $z\gtrsim 1$. Additionally we also see from the GWTC-3 [4] catalogue that the spin distributions of the remnants cannot still be accurately inferred, while the distributions about eccentricity remain unknown. We will again opt for the most simplistic formulation possible given the circumstances.
Consequently, we will work with both uniform as well as non-unform (Kroupa) IMFs [43] in our analysis. The uniform IMF, though probably unphysical still allows us to be maximally agnostic, while the Kroupa model is motivated by the stellar observations from the galactic plane. Also, we will continue to assume no effects of eccentricity in the binary orbit or BH spin in our analysis. Moreover, our evolution does not include other astrophysical effects, like Roche overflows, stellar winds or common envelope binary evolution. We remark that factoring in all of these components that we have left out require a fully numerical population synthesis, which is beyond the scope of this work. Therefore, what we are assuming is that the dynamical friction and energy loss by GWs are the only dissipative components in the dynamics. So, any constraint on the density of the ULDM from our analysis should be treated only as an upper limit.
Let us now focus on our choice of lower and upper cut-off for the IMFs. It is reasonable to expect that progenitor stars which have masses below a limit cannot produce compact remnants. Keeping this in mind we choose a lower cut-off of the progenitor masses given by $M_{i}^{\rm P}\geq 10M_{\odot}$. Then the use of the uniform IMF requires us to also choose an appropriate upper bound in the progenitor stellar-mass distribution. Masses of stars are bounded from above to $M_{i}^{\rm P}\leq 100M_{\odot}$ from considerations of hydrostatic equilibrium. For our purposes in this work, we choose a slightly lower value for the upper limit of the progenitor masses, namely, $M_{i}^{\rm P}\leq 80M_{\odot}$. A limitation of this mass cut-off could be that it might produce bias for the choice of uniform IMF. However, the bias for the case of the more realistic non-uniform IMF will be negligible, as that IMF function decays to zero at those high mass values.
Figure 3: Plot showing the composition of merged and unmerged final states of all possible combinations of binary main sequence star progenitors beginning at $z=0.5$. The grey shaded regions are off-limits in progenitor configuration in our modelling (in those regions, either masses are larger than $80M_{\odot}$, or smaller than $10M_{\odot}$). We show the $z$ distribution for the merged cases (green shades) and the binary separation $a$ in units of solar radius $\mathrm{R}_{\odot}$ at $z=0$ for the unmerged cases as a function of the total progenitor masses in the binary and assuming a mass transfer, for different cases of dark matter density. The red line is the separatrix between the unmerged and merged populations at $z=0$. See text for further details.
A final uncertainty which plays a role in determining the merger timescales is the choice of initial separation of the progenitor stars. However, there are still some physically motivated factors which help to narrow down the range of the initial separations. We note that the initial separation is bound from below if we demand no Roche overflow in the initial progenitor configuration. Consequently we assume the initial configuration of the binaries to always obey the no Roche overflow condition. We have implemented this by making the sum of the initial Roche lobes of the component stars $2.5$ times the sum of their radii. The rescaling of merger rate (see Sec  IV) ensures that the rate of mergers in our model is commensurate with observations even with our specific choice of initial separations. We now return to the question of the final state of main sequence star progenitors, all of which evolve in the presence of ULDM, while starting at a fixed redshift. We present the answer to this question in Fig. 3, where we show that the end states of the evolution for progenitor binaries starting from $z=0.5$. Here and henceforth, we adopt the convention that for the progenitors $M_{2}^{\rm P}\leq M_{1}^{\rm P}$, such that the mass-ratio $q^{\rm P}\leq 1$. Following which, in Fig. 3, we show four different panels correspond to four different ambient densities of the ULDM environment. For each panel, we show the mass ratio of the progenitors $q^{\rm P}$ on the x-axis and the total mass of the binary configuration $M_{\rm T}^{\rm P}=M_{1}^{\rm P}+M_{2}^{\rm P}$ on the y-axis. Following our reasoning above, the grey shaded regions represent configurations that are off-limits to our interest, i.e., with progenitors masses being too high, or, too low. It is expected that the merger timescales for the binary systems will not all be the same, and thus we would get a combination of merged and unmerged binary populations, depending on the DM density in the environment. In Fig. 3, the red line acts as a separatrix, which separates the configuration space into merged and unmerged endpoints for the binaries. The redshifts for the merged population are shown by the green-yellow shade, while for the unmerged populations we depict the binary separation at $z=0$ by the blue shade.
As evident from Fig. 3, an increase in the ULDM density leads to a reduction in the area under the blue-shaded region, which, as expected, implies more mergers occurring at higher redshifts. Additionally, we observe that a higher ULDM density facilitates the merging of remnants that would otherwise remain unmerged, particularly those originating from more massive progenitors. The oscillatory pattern seen in the parameter space of the unmerged population for $\rho=5\times 10^{7}~\textrm{GeV}/\textrm{cm}^{3}$ (top left plot of Fig. 3) arises from similar oscillatory non-linearities present in our assumed mass-transfer function. Moreover, as the ULDM density increases, its influence becomes more pronounced, eventually overshadowing the mass-transfer effects. This leads to a smoother and more distinct separation between merged and unmerged regions, as observed for higher values of $\rho$, e.g., $\rho=12.5\times 10^{7}~\textrm{GeV}/\textrm{cm}^{3}$ (bottom right plot of Fig. 3).
To further quantify these effects, we have presented the fraction of the merged configurations along with the ULDM densities within which they evolve in Table 1. Arising out of which we could identify a range of the density of the ULDM environment, namely $1.0\times 10^{7}\mathrm{GeV}/\textrm{cm}^{3}\leq\rho\leq 5.0\times 10^{7}\mathrm{GeV}/\textrm{cm}^{3}$, within which the ULDM remains too rare to induce mergers for binaries starting from $z=0.5$. However, for densities $\rho\gtrsim 5\times 10^{7}\mathrm{GeV}/\textrm{cm}^{3}$, ULDM begins to have a significant impact on the merger fraction, leading to nearly $80\%$ of configurations merging for densities $\sim 10^{8}\mathrm{GeV}/\textrm{cm}^{3}$ (see Table 1). These findings indicate that, within the framework of our simplified toy model, the presence of ULDM influences the evolution of the compact binary populations, systematically increasing the fraction of mergers occurring at earlier cosmic times. While this model is intentionally minimalistic, it serves as an illustrative tool to study how the surrounding dark matter environment can shape merger statistics.
ULDM density in $\mathrm{GeV}/\textrm{cm}^{3}$ Merger Fraction
$20\times 10^{7} 0.0$ %
$1.00\times 10^{7}$ 0.0 %
$5.00\times 10^{7}$ 43.86 %
$7.50\times 10^{7}$ 72.56 %
$10.0\times 10^{7}$ 78.89 %
$12.5\times 10^{7}$ 80.78 %
$15.0\times 10^{7}$ 81.26 %
Table 1: Table showing the fraction of progenitors that have merged at $z=0$, starting their orbital evolution at $z=0.5$, with an increase in the ULDM density $\rho$. Note that we use a value of $m_{\Phi}=10^{-15}\text{eV}$ to be consistent with the ultralight condition $M\mu\ll 1$. The conservation between the unit for the DM density used in this work with the density in $\textrm{gm}/\textrm{cm}^{3}$ is: $1~\text{GeV}/\textrm{cm}^{3}=1.78\times 10^{-24}~\text{gm}/\textrm{cm}^{3}$.
IV Effects on Astrophysical Merger Rates
The results of the previous section have firmly established that presence of an ULDM environment should have an effect on the population statistics of mergers. However, we only considered progenitor sources at a single redshift $z=0.5$. A complete picture can be obtained if we can repeat the above exercise, but factoring in sources from a given range of redshifts. Additionally, we recall that all redshifts will not contribute equally to the generation of compact binary progenitors. This is because the comoving density of main-sequence stars is not constant with the redshift, and has a functional dependence which is well-known in the literature as the Star Formation Rate (SFR). In our analysis in this section, we have used the SFR function, as proposed in Madau & Dickinson [45]. At this outset, we would like to emphasize that the SFR simply provides the total mass of the stars produced at a given redshift per comoving volume. One needs to convolve it with the IMF in order to get the number density of the stars. For the cases we have considered, the convolution integrals will be non-trivial for the Kroupa IMF and just a constant factor in the case of flat IMF. Subsequent multiplication with the comoving volume results into a function for the number of sources in the redshift bin $z$ and $z+dz$, which can be observed per unit observer’s time. This function, representing number of binary progenitors per unit redshift per unit time, is given by
$\frac{d^{2}N}{dzdt}=\frac{1}{1+z}\left(\frac{dV_{\rm c}}{dz}\right)\mathcal{R_{\psi}}(z),$ (15)
where $\mathcal{R_{\psi}}(z)$ is the SFR at redshift $z$, and $(dV_{\rm c}/dz)$ is arising from the cosmological comoving volume. The factor $(1+z)^{-1}$ appears because of the difference in the time intervals at the source and at the observer. Given Eq. (15), we are now in a position to compute the rate of binary compact mergers for our underlying astrophysical model. To compute the ‘merger-rate density’ we need to first evolve all sources at a given $z$, given by Eq. (15), either to their merger or to $z=0$. We then bin the population of all the merged events according to their merger redshifts, which gives for us the number $(d^{2}N_{\rm m}/dz_{\rm m}dt)$, depicting the merger events between redshifts $z_{\rm m}$ and $z_{\rm m}+dz_{\rm m}$, observed per unit observer’s time. Here, $z_{m}$ represents the merger redshifts. Dividing the above by the comoving volume factor at $z_{m}$ and multiplying by a factor of $1+z_{\rm m}$ to account for the redshift factor, we finally arrive at our estimate for the merger rate density,
$\mathcal{R}_{\rm m}(z_{\rm m})\sim\frac{d^{2}N_{\rm m}}{dz_{\rm m}dt}\left(\frac{dz_{\rm m}}{dV_{\rm c}}\right)\left(1+z_{\rm m}\right)~.$ (16)
We can now compare our results and match our estimates of the merger rate $\mathcal{R}_{\rm m}(z_{\rm m})$ to the inferred rate $\mathcal{R}_{\rm m}^{\mathrm{inf}}(z_{\rm m})$ from the LIGO-Virgo observations. But before we do that a word of caution is in order. We reiterate that our baseline model does not consider inputs from many dissipative astrophysical effects (like Roche overflows, common envelope evolutions etc.). It also does not contain information about other non-dissipative astrophysics like for example the $z$ dependence of the stellar IMF or the fraction of progenitors that get trapped in binary systems. Therefore, it is natural to expect an incorrect estimation for our calculated merger rates. We should also remark that at current sensitivity, GW observations can only observe the ‘local’ merger rate density $\mathcal{R}_{\rm m}^{\mathrm{obs}}(0)$, i.e., the merger rate density at $z=0$. The inaccuracy in the estimation of the merger rate can nevertheless be removed by an overall rescaling of the number of sources such that the rescaled estimates of the rate $\mathcal{R}_{\rm m}(z_{\rm m})$ obey observational bounds at $z_{\rm m}=0$. In other words, such a rescaling will make $\mathcal{R}_{\rm m}(0)$ lie within the error bars of $\mathcal{R}_{\rm m}^{\mathrm{obs}}(0)\equiv R_{0}$. Subsequently, following this standard procedure in the population synthesis literature (see, for example [17]) we make use of appropriate rescaling functions to bring the value of $\mathcal{R}_{\rm m}(0)$ within the GWTC-3 bounds [5].
Figure 4: Plot showing merger rate densities $\mathcal{R}_{m}(z_{m})$ in blue as a function of redshift $z$ for our choices of ULDM density $\rho$. The green shaded region represents our current uncertainty about the total local merger rate density $R_{0}$ of binary compact objects inferred from GWTC-3 data [5]. The red dashed line show the inferred behaviour of the exponent $\kappa$ of $\mathcal{R}_{m}^{\mathrm{inf}}(z_{m})$ from GWTC-3 data along with its error bars. See text for more details.
Then in order to make predictions for mergers at higher $z_{m}$ we have to extrapolate the trends that are observed within the sensitivity horizon of the respective detectors. This is usually done by a inferring a power-law behaviour given by a parameter $\kappa$ for higher $z_{\rm m}$ from GWTC-3 data. It also should be noted that the bounds on $\kappa$ come from the observations of BBH systems alone. This is simply because only BBH systems can be observed in enough numbers at $z_{\rm m}\lesssim 1$ to make inferences about $\kappa$ using the GWTC-3 detector sensitivity. The overall $z_{\rm m}$-dependent statistics of merging compact binaries can then be modelled by a function with two parameters $R_{0}$ and $\kappa$. We note that the values of these parameters are based on direct observation and extrapolations from inferences respectively. Proceeding this way, we finally arrive at the inferred merger rate $\mathcal{R}_{\rm m}^{\mathrm{inf}}(z_{\rm m})$ given by,
$\mathcal{R}_{\rm m}^{\mathrm{inf}}(z_{\rm m})\simeq R_{0}\left(1+z_{\rm m}\right)^{\kappa}~.$ (17)
The greatest uncertainty to $R_{0}$ (in units of $\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$) is contributed by the binary NSs, yielding a very wide range for the merger rate $10\leq R_{0}\leq 1700$ [5]. Consequently we have scaled our populations to have $R_{0}=500~\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, so that our results are consistent with the bounds on $R_{0}$ presented by GWTC-3.
Following the previous discussion, we have plotted our rescaled merger rate estimates (blue lines), and the median rate inferred from the GWTC-3 (red dashed lines) for four different choices of the ULDM density in Fig. 4. Then assuming the same allowed value of $R_{0}=500~\mathrm{Gpc}^{-3}\mathrm{yr}^{-1}$, we plot the uncertainty region mapped by the $90\%$ credible values of $\kappa$ as a red shade (see Fig. 4). The uncertainty in $R_{0}$ is also depicted as a green shaded region in Fig. 4. Following the convention in [8], the extent of $R_{0}$ along the redshift is modelled till $z\sim 0.1$, which is roughly the horizon size of the O3 LVK detectors. We also remark that the reconstruction of $\mathcal{R}_{m}^{\mathrm{inf}}(z_{m})$ from GWTC-3 is mainly based on low redshift ($z_{\rm m}\lesssim 1$) events, and is therefore not likely to be an accurate representation at $z_{\rm m}\gtrsim 2$.
The rate estimates plotted in Fig. 4 form our second set of results. We note that in spite of the rescaling applied to $\mathcal{R}_{\rm m}(z_{\rm m})$, its $z_{\rm m}\lesssim 1$ behaviour for ULDM density $\rho\leq 10^{4}\text{GeV}/\textrm{cm}^{3}$ does not respect the $90\%$ bounds on the power-law exponent as set by GWTC-3. This indicates that the corresponding models can be ruled out at $90\%$ confidence from the GWTC-3 data [^2] Of course, there are several astrophysical uncertainties in our baseline model, starting from constant IMF, to ignoring Roche overflow.. Note that we can rule out only the model, which is a combination of our baseline model plus ULDM, given by a density of $\rho\leq 10^{4}\text{GeV}/\textrm{cm}^{3}$. We cannot however constrain ULDM $\rho$ in an absolute sense, because the effect of ULDM can be degenerate with the effects of additional dissipative factors like Roche overflow, stellar winds or CE phase evolution. Proceeding in this way, we see that for $\rho=10^{8}\text{GeV}/\textrm{cm}^{3}$, our model is only marginally consistent with the inferred merger rate for $z_{\rm m}\sim 0.1$. Finally, ULDM model with DM density, given by $\rho=10^{12}\text{GeV}/\textrm{cm}^{3}$ is the best performing. First of all, we see that our merger rate estimation for ULDM density of $\rho=10^{12}\text{GeV}/\textrm{cm}^{3}$ agrees with the GWTC power-law bounds for $z_{\rm m}\lesssim 2$, after which the agreement is lost. This disagreement is actually considerably softened, given that the GWTC power-law bounds themselves are not reliable at such $z_{m}$.
The merger rate density also decides the merger probability, which is simply the probability of having a binary merger event between the redshifts of $z_{\rm m}$ and $z_{\rm m}+dz_{\rm m}$, given by the time integral of the quantity $(d^{2}N_{\rm m}/dz_{\rm m}dt)$ defined above. In other words,
$p_{\mathrm{merg}}(z_{\rm m})\sim\int dt\frac{d^{2}N_{\rm m}}{dz_{\rm m}dt}~.$ (18)
We show the merger probabilities of GW events as a function of the merger redshifts $z_{\rm m}$, for the present scenario, in Fig. 5. We note that our findings in Fig. 5 are consistent with our previous results. Specifically, we obtained in Section III.1 that denser ULDM environments tend to reduce merger timescales and make individual binaries marge faster, leading to a greater proportion of merged binaries among progenitors from a given source. In terms of a population with a source distribution, this simply means that with shrinking merger timescales, the peak of the resulting merger distributions should move towards higher values of $z_{\rm m}$. We see that this prediction is exactly reproduced by our results for $p_{\mathrm{merg}}$ in Fig. 5. While the ULDM density $\rho=10^{4}\text{GeV}/\textrm{cm}^{3}$ is indistinguishable from the baseline model without DM, for densities $\rho\geq 10^{8}\text{GeV}/\textrm{cm}^{3}$ the merger probability indeed shifts to the right. The redshift at the peak of the merger probability distribution is thus an observable, which carries the signature of ULDM density $\rho$. However, with the current GW detector sensitivity, it is not possible to faithfully recreate the merger probability distribution of the GW events. We remark that improvements in ground-based GW detectors would overcome this difficulty and put another independent constraint upon the ULDM environments through direct observation of the peak in the merger redshift for merger probability distribution.
Figure 5: Plot showing the merger probability distribution as a function of the redshift $z$ for four different choices of the ULDM density $\rho$. As evident, increasing $\rho$ leads to more merger happening at larger redshifts, leading to a shift of the peak of the merger distribution towards higher redshifts. See text for more details.
V Conclusion
Given that astrophysical environments have previously been found to have non-trivial effects on the GW phasing and merger timescale of individual binaries, it is natural to expect that these environments also leave a significant imprint on their merger statistics. However, baryonic environments are highly case-specific and cannot be assumed to be present for every member of a merger population. In contrast, this situation is notably different for ULDM, as DM is expected to be present in all galaxies, and ULDM are also possibly present in the core region of a galaxy to avoid the core-cusp problem, where compact binary mergers can potentially occur. This ubiquity makes ULDM an intriguing candidate for studying environmental effects on merger statistics. Specifically, additional energy loss through dynamical friction, as the binaries moves in a DM environment, is expected to influence the distribution of the merger timescales, systematically shifting the merger populations toward shorter timescales. It is within this framework that the significance of our work becomes apparent.
In this study, we have performed a series of investigations aimed at quantifying the effects of ULDM and examining whether and how it can influence merger statistics of binaries. Our approach employs a simplified implementation of both the baseline astrophysical model and the DM medium. This is, however, just a toy model, constructed with the primary objective of providing a minimal yet insightful framework to illustrate and analyze the influence of a dark matter environment on merger statistics. The goal is not to present a fully realistic astrophysical model but rather to explore the key qualitative effects that arise when compact binaries evolve within a ULDM background.
Despite these simplifying assumptions, our analysis clearly demonstrates that ULDM fields can indeed impact merger statistics, particularly when their density exceeds the threshold of $\rho>10^{4}\text{GeV}/\textrm{cm}^{3}$. We started by considering individual binaries and then probing how their merger time gets modified. As expected, and as clearly depicted in Fig. 1 and Fig. 2, the binaries in the DM environment merges quickly compared to the binary that evolves in vacuum. For example, in the case of a binary with total mass $45M_{\odot}$ and mass-ratio of $0.8$, the binary with a ULDM environment with DM density $\rho=10^{8}\text{GeV}/\textrm{cm}^{3}$ merges at $z=0.4$, while in absence of DM environment merges at $z=0.07$, while they both start at $z=0.8$. Thus there is a significant effect on the merger timescale.
This feature continues to hold even for a binary population, where also, the presence of ULDM leads to significantly higher proportion of mergers compared to the vacuum case. In particular, for $\rho>10^{8}\text{GeV}/\textrm{cm}^{3}$, almost $80\%$ of the original progenitor binary configurations have merged. This also influences the merger rate of binaries. As evident from Fig. 4, the DM densities below the threshold value of $10^{4}\text{GeV}/\textrm{cm}^{3}$, the merger rate is in tension with the LVK estimation of the merger rate. This seems to be a tantalizing evidence for the existence of DM, but then the uncertainties in the baseline model, along with various astrophysical considerations can tame this result. Thus, we would like to conclude that merger rate densities can provide a gold mine of information about the existence, and possibly, the nature of DM, provided we have better estimates of the merger rate at higher redshifts and more refined baseline model. Specifically, we have shown that increasing the density of the DM medium leads to a shift in the peak of the merger probability distribution towards higher redshifts. This is expected, as the DM density increases the dynamical friction also increases, leading to an enhanced loss of the orbital energy for the dipole. As a consequence, most of the binaries lose energy faster and hence merges at a higher redshift, i.e., earlier than expected from a vacuum environment. Thus the peak of the merger probability distribution in redshift is another observable that can tell us a lot about the existence and nature of the DM environment in galaxies. These results highlight the potential role of ULDM in shaping the observable characteristics of compact binary mergers and hence provides a novel perspective on how DM may influence astrophysical merger events.
Acknowledgements
The authors acknowledge the use of the Noether workstation at the Department of Physics, IIT Gandhinagar. The research of KC is supported by the PPLZ grant (Project No. 10005320/0501) of the Czech Academy of Sciences. The research of SS is supported by the Department of Science and Technology, Government of India, under the ANRF CRG Grant (No. CRG/2023/000545). The research of SC is supported by MATRICS (MTR/2023/000049) and Core Research (CRG/2023/000934) Grants from SERB, ANRF, Government of India. SC also thanks the local hospitality at ICTS through the associateship program, where a part of this work was done.